### v4

# 3D Human Mesh Estimation from Virtual Markers

Abstract
Inspired by the success of volumetric 3D pose estimation, some recent human mesh estimators propose to estimate 3D skeletons as intermediate representations, from which, the dense 3D meshes are regressed by exploiting the mesh topology. However, body shape information is lost in extracting skeletons, leading to mediocre performance. The advanced motion capture systems solve the problem by placing dense physical markers on the body surface, which allows to extract realistic meshes from their non-rigid motions. However, they cannot be applied to wild images without markers. In this work, we present an intermediate representation, named virtual markers, which learns 64 landmark keypoints on the body surface based on the large-scale mocap data in a generative style, mimicking the effects of physical markers. The virtual markers can be accurately detected from wild images and can reconstruct the intact meshes with realistic shapes by simple interpolation. Our approach outperforms the state-of-the-art methods on three datasets. In particular, it surpasses the existing methods by a notable margin on the SURREAL dataset, which has diverse body shapes. Code is available at https://github.com/ShirleyMaxx/VirtualMarker.
1 Introduction
Figure 1: Mesh estimation results on four examples with different body shapes. Pose2Mesh [7] which uses 3D skeletons as the intermediate representation fails to predict accurate shapes. Our virtual marker-based method obtains accurate estimates.
3D human mesh estimation aims to estimate the 3D positions of the mesh vertices that are on the body surface. The task has attracted a lot of attention from the computer vision and computer graphics communities [3, 43, 30, 35, 51, 19, 25, 37, 27, 10] because it can benefit many applications such as virtual reality [15]. Recently, the deep learning-based methods [19, 7, 29] have significantly advanced the accuracy on the benchmark datasets.
The pioneer methods [51, 19] propose to regress the pose and shape parameters of the mesh models such as SMPL [36] directly from images. While straightforward, their accuracy is usually lower than the state-of-the-arts. The first reason is that the mapping from the image features to the model parameters is highly non-linear and suffers from image-model misalignment [29]. Besides, existing mesh datasets [16, 54, 38, 28] are small and limited to simple laboratory environments due to the complex capturing process. The lack of sufficient training data severely limits its performance.
Recently, some works [26, 39] begin to formulate mesh estimation as a dense 3D keypoint detection task inspired by the success of volumetric pose estimation [47, 50, 65, 44, 59, 45]. For example, in [26, 39], the authors propose to regress the 3D positions of all vertices. However, it is computationally expensive because it has more than several thousand vertices. Moon and Lee [39] improve the efficiency by decomposing the 3D heatmaps into multiple 1D heatmaps at the cost of mediocre accuracy. Choi et al. [7] propose to first detect a sparser set of skeleton joints in the images, from which the dense 3D meshes are regressed by exploiting the mesh topology. The methods along this direction have attracted increasing attention [7, 29, 55] due to two reasons. First, the proxy task of 3D skeleton estimation can leverage the abundant 2D pose datasets which notably improves the accuracy. Second, mesh regression from the skeletons is efficient. However, important information about the body shapes is lost in extracting the 3D skeletons, which is largely overlooked previously. As a result, different types of body shapes, such as lean or obese, cannot be accurately estimated (see Figure 1).
The professional marker-based motion capture (mocap) method MoSh [35] places physical markers on the body surface and explore their subtle non-rigid motions to extract meshes with accurate shapes. However, the physical markers limit the approach to be used in laboratory environments. We are inspired to think whether we can identify a set of landmarks on the mesh as virtual markers, e.g., elbow and wrist, that can be detected from wild images, and allow to recover accurate body shapes? The desired virtual markers should satisfy several requirements. First, the number of markers should be much smaller than that of the mesh vertices so that we can use volumetric representations to efficiently estimate their 3D positions. Second, the markers should capture the mesh topology so that the intact mesh can be accurately regressed from them. Third, the virtual markers have distinguishable visual patterns so that they can be detected from images.
In this work, we present a learning algorithm based on archetypal analysis [12] to identify a subset of mesh vertices as the virtual markers that try to satisfy the above requirements to the best extent. Figure 2 shows that the learned virtual markers coarsely outline the body shape and pose which paves the way for estimating meshes with accurate shapes. Then we present a simple framework for 3D mesh estimation on top of the representation as shown in Figure 3. It first learns a 3D keypoint estimation network based on [47] to detect the 3D positions of the virtual markers. Then we recover the intact mesh simply by interpolating them. The interpolation weights are pre-trained in the representation learning step and will be adjusted by a light network based on the prediction confidences of the virtual markers for each image.
We extensively evaluate our approach on three benchmark datasets. It consistently outperforms the state-of-the-art methods on all of them. In particular, it achieves a significant gain on the SURREAL dataset [53] which has a variety of body shapes. Our ablation study also validates the advantages of the virtual marker representation in terms of recovering accurate shapes. Finally, the method shows decent generalization ability and generates visually appealing results for the wild images.
2 Related work
2.1 Optimization-based mesh estimation
Before deep learning dominates this field, 3D human mesh estimation [35, 2, 28, 41, 60] is mainly optimization-based, which optimizes the parameters of the human mesh models to match the observations. For example, Loper et al. [35] propose MoSh that optimizes the SMPL parameters to align the mesh with the 3D marker positions. It is usually used to get GT 3D meshes for benchmark datasets because of its high accuracy. Later works propose to optimize the model parameters or mesh vertices based on 2D image cues [2, 28, 41, 60, 11]. They extract intermediate representations such as 2D skeletons from the images and optimize the mesh model by minimizing the discrepancy between the model projection and the intermediate representations such as the 2D skeletons. These methods are usually sensitive to initialization and suffer from local optimum.
Figure 2: Left: The learned virtual markers (blue balls) in the back and front views. The grey balls mean they are invisible in the front view. The virtual markers act similarly to physical body markers and approximately outline the body shape. Right: Mesh estimation results by our approach, from left to right are input image, estimated 3D mesh overlayed on the image, and three different viewpoints showing the estimated 3D mesh with our intermediate predicted virtual markers (blue balls), respectively.
2.2 Learning-based mesh estimation
Recently, most works follow the learning-based framework and have achieved promising results. Deep networks [51, 19, 25, 37, 27] are used to regress the SMPL parameters from image features. However, learning the mapping from the image space to the parameter space is highly non-linear [39]. In addition, they suffer from the misalignment between the meshes and image pixels [62]. These problems make it difficult to learn an accurate yet generalizable model.
Some works propose to introduce proxy tasks to get intermediate representations first, hoping to alleviate the learning difficulty. In particular, intermediate representations of physical markers [61], IUV images [57, 62, 64, 63], body part segmentation masks [52, 24, 28, 40] and body skeletons [49, 7, 29, 55] have been proposed. In particular, THUNDR [61] first estimates the 3D locations of physical markers from images and then reconstructs the mesh from the 3D markers. The physical markers can be interpreted as a simplified representation of body shape and pose. Although it is very accurate, it cannot be applied to wild images without markers. In contrast, body skeleton is a popular human representation that can be robustly detected from wild images. Choi et al. [7] propose to first estimate the 3D skeletons, and then estimate the intact mesh from them. However, accurate body shapes are difficult to be recovered from the oversimplified 3D skeletons.
Our work belongs to the learning-based class and is related to works that use physical markers or skeletons as intermediate representations. But different from them, we propose a novel intermediate representation, named virtual markers, which is more expressive to reduce the ambiguity in pose and shape estimation than body skeletons and can be applied to wild images.
3 Method
In this section, we describe the details of our approach. First, Section 3.1 introduces how we learn the virtual marker representation from mocap data. Then we present the overall framework for mesh estimation from an image in Section 3.2. At last, Section 3.3 discusses the loss functions and training details.
3.1 The virtual marker representation
We represent a mesh by a vector of vertex positions $\mathbf{x}\in\mathbb{R}^{3M}$ where $M$ is the number of mesh vertices. Denote a mocap dataset such as [16] with $N$ meshes as $\overset{\frown}{\mathbf{X}}=[\mathbf{x}_{1},\,...,\,\mathbf{x}_{N}]\in\mathbb{R}^{3M\times N}$. To unveil the latent structure among vertices, we reshape it to $\mathbf{X}\in\mathbb{R}^{3N\times M}$ with each column $\mathbf{x}_{i}\in\mathbb{R}^{3N}$ representing all possible positions of the $i^{\text{th}}$ vertex in the dataset [16].
The rank of $\mathbf{X}$ is smaller than $M$ because the mesh representation is smooth and redundant where some vertices can be accurately reconstructed by the others. While it seems natural to apply PCA [18] to $\mathbf{X}$ to compute the eigenvectors as virtual markers for reconstructing others, there is no guarantee that the virtual markers correspond to the mesh vertices, making them difficult to be detected from images. Instead, we aim to learn $K$ virtual markers $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{K}]\in\mathbb{R}^{3N\times K}$ that try to satisfy the following two requirements to the greatest extent. First, they can accurately reconstruct the intact mesh $\mathbf{X}$ by their linear combinations: $\mathbf{X}=\mathbf{Z}\mathbf{A}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is a coefficient matrix that encodes the spatial relationship between the virtual markers and the mesh vertices. Second, they should have distinguishable visual patterns in images so that they can be easily detected from images. Ideally, they can be on the body surface as the meshes.
We apply archetypal analysis [12, 4] to learn $\mathbf{Z}$ by minimizing a reconstruction error with two additional constraints: (1) each vertex $\mathbf{x}_{i}$ can be reconstructed by convex combinations of $\mathbf{Z}$, and (2) each marker $\mathbf{z}_{i}$ should be convex combinations of the mesh vertices $\mathbf{X}$:
[…]
where $\mathbf{A}=[\bm{\alpha}_{1},...,\bm{\alpha}_{M}]\in\mathbb{R}^{K\times M}$, each $\bm{\alpha}$ resides in the simplex ${\Delta}_{K}\triangleq\{\bm{\alpha}\in\mathbb{R}^{K}\,\mathrm{s.t.}\,\bm{\alpha}\succeq 0\,\text{and}\,{||\bm{\alpha}||}_{1}=1\}$, and $\mathbf{B}=[\bm{\beta}_{1},...,\bm{\beta}_{K}]\in\mathbb{R}^{M\times K}$, $\bm{\beta}_{j}\in{\Delta}_{M}$. We adopt Active-set algorithm [4] to solve objective (1) and obtain the learned virtual markers $\mathbf{Z}=\mathbf{X}\mathbf{B}\in\mathbb{R}^{3N\times K}$. As shown in [12, 4], the two constraints encourage the virtual markers $\mathbf{Z}$ to unveil the latent structure among vertices, therefore they learn to be close to the extreme points of the mesh and located on the body surface as much as possible.
Figure 3: Overview of our framework. Given an input image $\mathbf{I}$, it first estimates the 3D positions $\hat{\mathbf{P}}$ of the virtual markers. Then we update the coefficient matrix $\hat{\mathbf{A}}$ based on the estimation confidence scores $\mathbf{C}$ of the virtual markers. Finally, the complete human mesh can be simply recovered by linear multiplication $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
Type Formula Reconst. Error (mm) $\downarrow$
Original $||\mathbf{X}-\mathbf{X}\mathbf{B}\mathbf{A}||^{2}_{F}$ 11.67
Symmetric $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ 10.98
Table 1: The reconstruction errors using the original and the symmetric sets of markers on the H3.6M dataset [16], respectively. The errors are small indicating that they are sufficiently expressive and can reconstruct all vertices accurately.
Post-processing. Since human body is left-right symmetric, we adjust $\mathbf{Z}$ to reflect the property. We first replace each $\mathbf{z}_{i}\in\mathbf{Z}$ by its nearest vertex on the mesh and obtain $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. This step allows us to compute the left or right counterpart of each marker. Then we replace the markers in the right body with the symmetric vertices in the left body and obtain the symmetric markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$. Finally we update $\mathbf{B}$ and $\mathbf{A}$ by minimizing $||\mathbf{X}-\mathbf{X}\widetilde{\mathbf{B}}^{sym}\widetilde{\mathbf{A}}^{sym}||^{2}_{F}$ subject to $\widetilde{\mathbf{Z}}^{sym}=\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. More details are elaborated in the supplementary.
Figure 2 shows the virtual markers learned on the mocap dataset [16] after post-processing. They are similar to the physical markers and approximately outline the body shape which agrees with our expectations. They are roughly evenly distributed on the surface of the body, and some of them are located close to the body keypoints, which have distinguishable visual patterns to be accurately detected. Table 1 shows the reconstruction errors of using original markers $\mathbf{X}\mathbf{B}$ and the symmetric markers $\mathbf{X}\widetilde{\mathbf{B}}^{sym}$. Both can reconstruct meshes accurately.
3.2 Mesh estimation framework
On top of the virtual markers, we present a simple yet effective framework for end-to-end 3D human mesh estimation from a single image. As shown in Figure 3, it consists of two branches. The first branch uses a volumetric CNN [47] to estimate the 3D positions $\hat{\mathbf{P}}$ of the markers, and the second branch reconstructs the full mesh $\hat{\mathbf{M}}$ by predicting a coefficient matrix $\hat{\mathbf{A}}$:
$\small\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}.$ (2)
We will describe the two branches in more detail.
3D marker estimation. We train a neural network to estimate a 3D heatmap $\hat{\mathbf{H}}=[\hat{\mathbf{H}}_{1},\,...,\,\hat{\mathbf{H}}_{K}]\in\mathbb{R}^{K\times D\times H\times W}$ from an image. The heatmap encodes per-voxel likelihood of each marker. There are $D\times H\times W$ voxels in total which are used to discretize the 3D space. The 3D position $\hat{\mathbf{P}}_{z}\in\mathbb{R}^{3}$ of each marker is computed as the center of mass of the corresponding heatmap $\hat{\mathbf{H}}_{z}$ [47] as follows:
$\small\hat{\mathbf{P}}_{z}=\sum_{d=1}^{D}\sum_{h=1}^{H}\sum_{w=1}^{W}(d,h,w)\cdot\hat{\mathbf{H}}_{z}(d,h,w).$ (3)
The positions of all markers are represented as $\hat{\mathbf{P}}=[\hat{\mathbf{P}}_{1},\hat{\mathbf{P}}_{2},\cdots,\hat{\mathbf{P}}_{K}]$.
Interpolation. Ideally, if we have accurate estimates for all virtual markers $\hat{\mathbf{P}}$, then we can recover the complete mesh by simply multiplying $\hat{\mathbf{P}}$ with a fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ with sufficient accuracy as validated in Table 1. However, in practice, some markers may have large estimation errors because they may be occluded in the monocular setting. Note that this happens frequently. For example, the markers in the back will be occluded when a person is facing the camera. As a result, inaccurate markers positions may bring large errors to the final mesh if we directly multiply them with the fixed matrix $\widetilde{\mathbf{A}}^{sym}$.
Our solution is to rely more on those accurately detected markers. To that end, we propose to update the coefficient matrix based on the estimation confidence scores of the markers. In practice, we simply take the heatmap score at the estimated positions of each marker, i.e. $\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z})$, and feed them to a single fully-connected layer to obtain the coefficient matrix $\hat{\mathbf{A}}$. Then the mesh is reconstructed by $\hat{\mathbf{M}}=\hat{\mathbf{P}}\hat{\mathbf{A}}$.
3.3 Training
We train the whole network end-to-end in a supervised way. The overall loss function is defined as:
$\mathcal{L}=\lambda_{vm}\mathcal{L}_{vm}+\lambda_{c}\mathcal{L}_{conf}+\lambda_{m}\mathcal{L}_{mesh}.$ (4)
Virtual marker loss. We define $\mathcal{L}_{vm}$ as the $L_{1}$ distance between the predicted 3D virtual markers $\hat{\mathbf{P}}$ and the GT $\hat{\mathbf{P}}^{*}$ as follows:
$\mathcal{L}_{vm}=\|\hat{\mathbf{P}}-\hat{\mathbf{P}}^{*}\|_{1}.$ (5)
Note that it is easy to get GT markers $\hat{\mathbf{P}}^{*}$ from GT meshes as stated in Section 3.1 without additional manual annotations.
Confidence loss. We also require that the 3D heatmaps have reasonable shapes, therefore, the heatmap score at the voxel containing the GT marker position $\hat{\mathbf{P}}_{z}^{*}$ should have the maximum value as in the previous work [17]:
$\mathcal{L}_{conf}=-\sum_{z=1}^{K}log(\hat{\mathbf{H}}_{z}(\hat{\mathbf{P}}_{z}^{*})).$ (6)
Mesh loss. Following [39], we define $\mathcal{L}_{mesh}$ as a weighted sum of four losses:
$\small\mathcal{L}_{mesh}=\mathcal{L}_{vertex}+\mathcal{L}_{pose}+\mathcal{L}_{normal}+\lambda_{e}\mathcal{L}_{edge}.$ (7)
  • Vertex coordinate loss. We adopt $L_{1}$ loss between predicted 3D mesh coordinates $\hat{\mathbf{M}}$ with GT mesh $\hat{\mathbf{M}}^{*}$ as:
    $\mathcal{L}_{vertex}=\|\hat{\mathbf{M}}-\hat{\mathbf{M}}^{*}\|_{1}.$ (8)
  • Pose loss. We use $L_{1}$ loss between the 3D landmark joints regressed from mesh $\hat{\mathbf{M}}\mathcal{J}$ and the GT joints $\hat{\mathbf{J}}^{*}$ as:
    $\mathcal{L}_{pose}=\|\hat{\mathbf{M}}\mathcal{J}-\hat{\mathbf{J}}^{*}\|_{1},$ (9)
    where $\mathcal{J}\in\mathbb{R}^{M\times J}$ is a pre-defined joint regression matrix in SMPL model [2].
  • Surface losses. To improve surface smoothness [56], we supervise the normal vector of a triangle face with GT normal vectors by $\mathcal{L}_{normal}$ and the edge length of the predicted mesh with GT length by $\mathcal{L}_{edge}$:
    $\mathcal{L}_{normal}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\left<\frac{\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}}{\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}},\hat{\mathbf{n}}_{f}^{*}\right>\right|,$ (10)
    $\mathcal{L}_{edge}=\sum_{f}\sum_{\{i,j\}\subset f}\left|\|\hat{\mathbf{M}}_{i}-\hat{\mathbf{M}}_{j}\|_{2}-\|\hat{\mathbf{M}}_{i}^{*}-\hat{\mathbf{M}}_{j}^{*}\|_{2}\right|.$
    where $f$ and $\hat{\mathbf{n}}_{f}^{*}$ denote a triangle face in the mesh and its GT unit normal vector, respectively. $\hat{\mathbf{M}}_{i}$ denote the $i^{th}$ vertex of $\hat{\mathbf{M}}$. denotes GT.
4 Experiments
Method Intermediate H3.6M 3DPW
[…]
Arnab et al. [1] CVPR’19 2D skeleton - 77.8 54.3 - - 72.2
HMMR [20] CVPR’19 - - - 56.9 139.3 116.5 72.6
DSD-SATN [49] ICCV’19 3D skeleton - 59.1 42.4 - - 69.5
VIBE [23] CVPR’20 - - 65.9 41.5 99.1 82.9 51.9
TCMR [6] CVPR’21 - - 62.3 41.1 102.9 86.5 52.7
MAED [55] ICCV’21 3D skeleton - 56.3 38.7 92.6 79.1 45.7
SMPLify [2] ECCV’16 2D skeleton - - 82.3 - - -
HMR [19] CVPR’18 - 96.1 88.0 56.8 152.7 130.0 81.3
GraphCMR [26] CVPR’19 3D vertices - - 50.1 - - 70.2
SPIN [25] ICCV’19 - - - 41.1 116.4 96.9 59.2
DenseRac [57] ICCV’19 IUV image - 76.8 48.0 - - -
DecoMR [62] CVPR’20 IUV image - 60.6 39.3 - - -
ExPose [9] ECCV’20 - - - - - 93.4 60.7
Pose2Mesh [7] ECCV’20 3D skeleton 85.3 64.9 46.3 106.3 88.9 58.3
I2L-MeshNet [39] ECCV’20 3D vertices 65.1 55.7 41.1 110.1 93.2 57.7
PC-HMR [37] AAAI’21 3D skeleton - - - 108.6 87.8 66.9
HybrIK [29] CVPR’21 3D skeleton 65.7 54.4 34.5 86.5 74.1 45.0
METRO [32] CVPR’21 3D vertices - 54.0 36.7 88.2 77.1 47.9
ROMP [48] ICCV’21 - - - - 108.3 91.3 54.9
Mesh Graphormer[33] ICCV’21 3D vertices - 51.2 34.5 87.7 74.7 45.6
PARE [24] ICCV’21 Segmentation - - - 88.6 74.5 46.5
THUNDR [61] ICCV’21 3D markers - 55.0 39.8 88.0 74.8 51.5
PyMaf [64] ICCV’21 IUV image - 57.7 40.5 110.1 92.8 58.9
ProHMR [27] ICCV’21 - - - 41.2 - - 59.8
OCHMR [21] CVPR’22 2D heatmap - - - 107.1 89.7 58.3
3DCrowdNet [8] CVPR’22 3D skeleton - - - 98.3 81.7 51.5
CLIFF [31] ECCV’22 - - 47.1 32.7 81.2 69.0 43.0
FastMETRO [5] ECCV’22 3D vertices - 52.2 33.7 84.1 73.5 44.6
VisDB [58] ECCV’22 3D vertices - 51.0 34.5 85.5 73.5 44.9
Ours Virtual marker 58.0 47.3 32.0 77.9 67.5 41.3
Table 2: Comparison to the state-of-the-arts on H3.6M [16] and 3DPW [54] datasets. means using temporal cues. The methods are not strictly comparable because they may have different backbones and training datasets. We provide the numbers only to show proof-of-concept results.
4.1 Datasets and metrics
H3.6M [16]. We use (S1, S5, S6, S7, S8) for training and (S9, S11) for testing. As in [19, 7, 32, 33], we report MPJPE and PA-MPJPE for poses that are derived from the estimated meshes. We also report Mean Per Vertex Error (MPVE) for the whole mesh.
3DPW [54] is collected in natural scenes. Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used.
SURREAL [53] is a large-scale synthetic dataset with GT SMPL annotations and has diverse samples in terms of body shapes, backgrounds, etc. We use its training set to train a model and evaluate the test split following [7].
4.2 Implementation Details
We learn $64$ virtual markers on the H3.6M [16] training set. We use the same set of markers for all datasets instead of learning a separate set on each dataset. Following [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using MPI-INF-3DHP [38], UP-3D [28], and COCO [34] training set for experiments on the H3.6M and 3DPW datasets. We adapt a 3D pose estimator [47] with HRNet-W48 [46] as the image feature backbone for estimating the 3D virtual markers. We set the number of voxels in each dimension to be $64$, i.e. $D=H=W=64$ for 3D heatmaps. Following [19, 26, 39], we crop every single human region from the input image and resize it to $256\times 256$. We use Adam [22] optimizer to train the whole framework for $40$ epochs with a batch size of $32$. The learning rates for the two branches are set to $5\times 10^{-4}$ and $1\times 10^{-3}$, respectively, which are decreased by half after the $30^{th}$ epoch. Please refer to the supplementary for more details.
[…]
Representation
HMR [19] CVPR’18 - 85.1 73.6 55.4
BodyNet [52] ECCV’18 Skel. + Seg. 65.8 - -
GraphCMR [26] CVPR’19 3D vertices 103.2 87.4 63.2
SPIN [25] ICCV’19 - 82.3 66.7 43.7
DecoMR [62] CVPR’20 IUV image 68.9 52.0 43.0
Pose2Mesh [7] ECCV’20 3D skeleton 68.8 56.6 39.6
PC-HMR [37] AAAI’21 3D skeleton 59.8 51.7 37.9
DynaBOA [14] TPAMI’22 - 70.7 55.2 34.0
Ours Virtual marker 44.7 36.9 28.9
Table 3: Comparison to the state-of-the-arts on SURREAL [53] dataset. means training on the test split with 2D supervisions. “Skel. + Seg.” means using skeleton and segmentation together.
4.3 Comparison to the State-of-the-arts
Results on H3.6M. Table 2 compares our approach to the state-of-the-art methods on the H3.6M dataset. Our method achieves competitive or superior performance. In particular, it outperforms the methods that use skeletons (Pose2Mesh [7], DSD-SATN [49]), body markers (THUNDR) [61], or IUV image [62, 64] as proxy representations, demonstrating the effectiveness of the virtual marker representation.
Results on 3DPW. We compare our method to the state-of-the-art methods on the 3DPW dataset in Table 2. Our approach achieves state-of-the-art results among all the methods, validating the advantages of the virtual marker representation over the skeleton representation used in Pose2Mesh [7], DSD-SATN [49], and other representations like IUV image used in PyMAF [64]. In particular, our approach outperforms I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33] by a notable margin, which suggests that virtual markers are more suitable and effective representations than detecting all vertices directly as most of them are not discriminative enough to be accurately detected.
Results on SURREAL. This dataset has more diverse samples in terms of body shapes. The results are shown in Table 3. Our approach outperforms the state-of-the-art methods by a notable margin, especially in terms of MPVE. Figure 1 shows some challenging cases without cherry-picking. The skeleton representation loses the body shape information so the method [7] can only recover mean shapes. In contrast, our approach generates much more accurate mesh estimation results.
[…]
Representation H3.6M SURREAL
(a) Skeleton 64.4 53.6
(b) Rand virtual marker 63.0 50.1
(c) Virtual marker 58.0 44.7
Table 4: Ablation study of the virtual marker representation for our approach on H3.6M and SURREAL datasets. “Skeleton” means the sparse landmark joint representation is used. “Rand virtual marker” means the virtual markers are randomly selected from all the vertices without learning. (c) is our method, where the learned virtual markers are used.
4.4 Ablation study
Virtual marker representation. We compare our method to two baselines in Table 4. First, in baseline (a), we replace the virtual markers of our method with the skeleton representation. The rest are kept the same as ours (c). Our method achieves a much lower MPVE than the baseline (a), demonstrating that the virtual markers help to estimate body shapes more accurately than the skeletons. In baseline (b), we randomly sample $64$ from the $6890$ mesh vertices as virtual markers. We repeat the experiment five times and report the average number. We can see that the result is worse than ours, which is because the randomly selected vertices may not be expressive to reconstruct the other vertices or can not be accurately detected from images as they lack distinguishable visual patterns. The results validate the effectiveness of our learning strategy.
Figure 4: Mesh estimation results of different methods on H3.6M test set. Our method with virtual marker representation gets better shape estimation results than Pose2Mesh which uses skeleton representation. Note the waistline of the body and the thickness of the arm.
Figure 1 shows some qualitative results on the SURREAL test set. The meshes estimated by the baseline which uses skeleton representation, i.e. Pose2Mesh [7], have inaccurate body shapes. This is reasonable because the skeleton is oversimplified and has very limited capability to recover shapes. Instead, it implicitly learns a mean shape for the whole training dataset. In contrast, the mesh estimated by using virtual markers has much better quality due to its strong representation power and therefore can handle different body shapes elegantly. Figure 4 also shows some qualitative results on the H3.6M test set. For clarity, we draw the intermediate representation (blue balls) in it as well.
Figure 5: Visualization of the learned virtual markers of different numbers of $K=16,32,96$, from left to right, respectively.
Number of virtual markers. We evaluate how the number of virtual markers affects estimation quality on H3.6M [16] dataset. Figure 5 visualizes the learned virtual markers, which are all located on the body surface and close to the extreme points of the mesh. This is expected as mentioned in Section 3.1. Table 5 (GT) shows the mesh reconstruction results when we have GT 3D positions of the virtual markers in objective (1). When we increase the number of virtual markers, both mesh reconstruction error (MPVE) and the regressed landmark joint error (MPJPE) steadily decrease. This is expected because using more virtual markers improves the representation power. However, using more virtual markers cannot guarantee smaller estimation errors when we need to estimate the virtual marker positions from images as in our method. This is because the additional virtual markers may have large estimation errors which affect the mesh estimation result. The results are shown in Table 5 (Det). Increasing the number of virtual markers $K$ steadily reduces the MPVE errors when $K$ is smaller than $96$. However, if we keep increasing $K$, the error begins to increase. This is mainly because some of the newly introduced virtual markers are difficult to detect from images and therefore bring errors to mesh estimation.
$K$ GT Det
[…]
16 46.8 39.8 58.7 47.8
32 20.1 14.2 58.2 48.3
64 11.0 7.5 58.0 47.3
96 9.9 5.6 59.6 48.2
Table 5: Ablation study of the different number of virtual markers ($K$) on H3.6M [16] dataset. (GT) Mesh reconstruction results when GT 3D positions of the virtual markers are used in objective (1). (Det) Mesh estimation results obtained by our proposed framework when we use different numbers of virtual markers ($K$).
Figure 6: Mesh estimation comparison results when using (a) fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$, and (b) updated $\hat{\mathbf{A}}$. Please zoom in to better see the details.
Figure 7: Top: Meshes estimated by our approach on images from 3DPW test set. The rightmost case in the dashed box shows a typical failure. Bottom: Meshes estimated by our approach on Internet images with challenging cases (extreme shapes or in a long dress).
Coefficient matrix. We compare our method to a baseline which uses the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$. We show the quality comparison in Figure 6. We can see that the estimated mesh by a fixed coefficient matrix (a) has mostly correct pose and shape but there are also some artifacts on the mesh while using the updated coefficient matrix (b) can get better mesh estimation results. As shown in Table 6, using a fixed coefficient matrix gets larger MPVE and MPJPE errors than using the updated coefficient matrix. This is caused by the estimation errors of virtual markers when occlusion happens, which is inevitable since the virtual markers on the back will be self-occluded by the front body. As a result, inaccurate marker positions would bring large errors to the final mesh estimates if we directly use the fixed matrix.
[…]
(a) Ours (fixed) 64.7 51.6
(b) Ours 58.0 47.3
Table 6: Ablation study of the coefficient matrix for our approach on H3.6M dataset. “fixed” means using the fixed coefficient matrix $\widetilde{\mathbf{A}}^{sym}$ to reconstruct the mesh.
4.5 Qualitative Results
Figure 7 (top) presents some meshes estimated by our approach on natural images from the 3DPW test set. The rightmost case shows a typical failure where our method has a wrong pose estimate of the left leg due to heavy occlusion. We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates. We further analyze how inaccurate virtual markers would affect the mesh estimation, i.e. when part of human body is occluded or truncated. According to the finally learned coefficient matrix $\mathbf{\hat{A}}$ of our model, we highlight the relationship weights among virtual markers and all vertices in Figure 8. We can see that our model actually learns local and sparse dependency between each vertex and the virtual markers, e.g. for each vertex, the virtual markers that contribute the most are in a near range as shown in Figure 8 (b). Therefore, in inference, if a virtual marker has inaccurate position estimation due to occlusion or truncation, the dependent vertices may have inaccurate estimates, while the rest will be barely affected. Figure 2 (right) shows more examples where occlusion or truncation occurs, and our method can still get accurate or reasonable estimates robustly. Note that when truncation occurs, our method still guesses the positions of the truncated virtual markers.
Figure 8: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{\hat{A}}$. (b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most. We can see that the dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
Figure 7 (bottom) shows our estimated meshes on challenging cases, which indicates the strong generalization ability of our model on diverse postures and actions in natural scenes. Please refer to the supplementary for more quality results. Note that since the datasets do not provide supervision of head orientation, face expression, hands, or feet, the estimates of these parts are just in canonical poses inevitably. Apart from that, most errors are due to inaccurate 3D virtual marker estimation which may be addressed using more powerful estimators or more diverse training datasets in the future.
5 Conclusion
In this paper, we present a novel intermediate representation Virtual Marker, which is more expressive than the prevailing skeleton representation and more accessible than physical markers. It can reconstruct 3D meshes more accurately and efficiently, especially in handling diverse body shapes. Besides, the coefficient matrix in the virtual marker representation encodes spatial relationships among mesh vertices which allows the method to implicitly explore structure priors of human body. It achieves better mesh estimation results than the state-of-the-art methods and shows advanced generalization potential in spite of its simplicity.
Acknowledgement
This work was supported by MOST-2022ZD0114900 and NSFC-62061136001.
Appendix
We elaborate on the post-processing implementation of the virtual markers and provide additional experimental details and results. At last, we discuss data from human subjects and the potential societal impact.
A.  Post-processing on Virtual Markers
As described in Section 3.1, considering the left-right symmetric human body structure, we slightly adjust the learned virtual markers $\mathbf{Z}$ to be symmetric. In fact, after the first step that updates each $\mathbf{z}_{i}$ by its nearest vertex to get $\widetilde{\mathbf{Z}}\in\mathbb{R}^{3\times K}$. $\widetilde{\mathbf{Z}}$ are almost symmetric with few exceptions. To get the final symmetric virtual markers $\widetilde{\mathbf{Z}}^{sym}\in\mathbb{R}^{3\times K}$, for each virtual marker located in the left body part, we take its symmetric vertex in the right body to be its symmetric counterpart.
Since the human mesh (i.e. SMPL [36]) itself is not strictly symmetric, we clarify the symmetric vertex pair (e.g. left elbow and right elbow) on a human mesh template $\mathbf{X}^{t}\in\mathbb{R}^{3\times M}$ in Figure 9. We place $\mathbf{X}^{t}$ at the origin of the 3D coordinate system. Formally, we define the cost of matching $i^{th}$ vertex to $j^{th}$ vertex to be $\bm{C}_{i,j}=\left|x_{i}+x_{j}\right|+\left|y_{i}-y_{j}\right|+\left|z_{i}-z_{j}\right|$. A symmetric vertex pair $(\mathbf{X}^{t}_{i},\mathbf{X}^{t}_{j})$ is defined to have the minimal cost $\bm{C}_{i,j}$. In this way, for each virtual marker in the left body, we take its symmetric vertex counterpart to be its symmetric virtual marker and finally get $\widetilde{\mathbf{Z}}^{sym}$.
Figure 9: Illustration of the human mesh template $\mathbf{X}^{t}$ at the 3D coordinate system and a symmetric vertex pair $(\mathbf{X}^{t}_{i},\mathbf{X}^{t}_{j})$.
B.  Experiments
In this section, we first add detailed descriptions for datasets and then provide more experimental results of our approach.
B.1  Datasets
H3.6M [16].
Following previous works [19, 25, 26, 42], we use the SMPL parameters generated from MoSh [35], which are fitted to the 3D physical marker locations, to get the GT 3D mesh supervision. Following standard practice [19], we evaluate the quality of 3D pose of $14$ joints derived from the estimated mesh, i.e. $\hat{\mathbf{M}}\mathcal{J}$. We report Mean Per Joint Position Error (MPJPE) and PA-MPJPE in millimeters (mm). The latter uses Procrustes algorithm [13] to align the estimates to GT poses before computing MPJPE. To evaluate mesh estimation results, we also report Mean Per Vertex Error (MPVE) which can be interpreted as MPJPE computed over the whole mesh.
3DPW [54].
The 3D GT SMPL parameters are obtained by using the data from IMUs when collected. Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set.
MPI-INF-3DHP [38]
is a 3D pose dataset with 3D GT pose annotations. Since this dataset does not provide 3D mesh annotations, following [19, 25], we only enforce supervision on the 3D skeletons (Eq. (9)) in mesh losses.
UP-3D [28]
is a wild 2D pose dataset with natural images. The 3D poses and meshes are obtained by SMPLify [2]. Due to the lack of GT 3D poses, the fitted meshes are not accurate. Therefore we only use the 2D annotations to train the 3D virtual marker estimation network as in [47].
COCO [34]
is a large wild 2D pose dataset with natural images. Previous work [39] used SMPLify-X [41] to obtain pseudo SMPL mesh annotations but they are not accurate. However, we find that if we project the 3D mesh to 2D image, the resulting 2D mesh vertices align well with the image. So we leverage the 2D annotations to train the virtual marker estimation network as in [47].
SURREAL [53]
is a large-scale synthetic dataset containing 6 million frames of synthetic humans. The images are photo-realistic renderings of people under large variations in shape, texture, viewpoint, and body pose. To ensure realism, the synthetic bodies are created using the SMPL body model, whose parameters are fit by the MoSh [35] given raw 3D physical marker data. All the images have a resolution of $320\times 240$. We use the same training split to train the model and evaluate the test split following [7].
B.2  Implementation Details and Computation Resource
Following common practice [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using the above 2D and 3D datasets for experiments on the H3.6M and 3DPW datasets. To leverage the 3D pose estimation dataset, i.e. MPI-INF-3DHP [38], we extend the $64$ virtual markers with the $17$ landmark joints (i.e. skeleton) from the MPI-INF-3DHP dataset. For experiments on the SURREAL dataset, we use its training set alone as in [7, 37]. We implement the proposed method with PyTorch. All the experiments are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs. The whole network is trained for 40 epochs with batch size 32 using Adam [22] optimizer.
We evaluate the model complexity in terms of FLOPs (G) and the number of model parameters in Table 7. Compared to the most recent state-of-the-art methods that directly regress all mesh vertices, such as I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33], our approach with virtual marker representation reduces the computation overhead by a large margin while getting better estimation quality. The last column shows the MPVE errors on 3DPW test set for performance reference.
[…]
I2L-MeshNet [39] ECCV’20 28.7 141.2 110.1
METRO [32] CVPR’21 153.0 397.5 88.2
Mesh Graphormer [33] ICCV’21 48.8 180.6 87.7
Ours 22.1 109.6 77.9
Table 7: Computation overhead comparison with the recent state-of-the-art methods that directly regress all 3D vertices. The rightmost column shows the MPVE errors on the 3DPW test set for performance reference.
Ours w/o $\mathcal{L}_{conf}$ w/o $\mathcal{L}_{pose}$ w/o $\mathcal{L}_{normal}$ w/o $\mathcal{L}_{edge}$
[…]
Table 8: MPVE errors on H3.6M [16] test set when ablating different loss terms.
[…]
None (Ours) 77.9 67.5 41.3
2 Arms 79.2 $\uparrow$ 1.3 68.2 $\uparrow$ 0.7 42.2 $\uparrow$ 0.9
2 Legs 78.3 $\uparrow$ 0.4 67.9 $\uparrow$ 0.4 41.7 $\uparrow$ 0.4
Body 78.6 $\uparrow$ 0.7 68.0 $\uparrow$ 0.5 41.8 $\uparrow$ 0.5
Random 78.7 $\uparrow$ 0.8 68.1 $\uparrow$ 0.6 41.9 $\uparrow$ 0.6
Table 9: Results on 3DPW [54] test set when different parts of virtual markers (VM) are occluded.
B.3  Additional Quantitative Results
Figure 10: Meshes estimated by our approach on Internet images with challenging cases (complex poses or extreme body shapes).
Different loss terms.
Table 8 reports the MPVE error on H3.6M [16] test set when ablating different loss terms. The confidence loss [17] is used to encourage the interpretability of the heatmaps to have a maxima response at the GT position. Without the confidence loss, the error increases slightly. If ablating the surface losses, MPVE increases a lot, which demonstrates the smoothing effect of these two terms.
Robustness to occlusion.
We report results when different virtual markers are occluded by a synthetic mask in Table 9. The errors are slightly larger than the original image (None), which validates the effectiveness of the locality of the virtual marker representation. Occluding arm regions results in a larger error increase. This may be because the arm has larger variations in the dataset.
Comparison to fitting.
In order to disentangle the ability of mesh regression from markers using $\hat{\mathbf{A}}$ with the ability to detect the virtual markers accurately from images, we first compute the estimation errors of the virtual markers. The MPJPE over all the virtual markers is $35.5$mm, which demonstrates that these virtual markers can be accurately detected from the images. We then fit the mesh model parameters to these virtual markers. Table 10 shows the metrics of the fitted mesh on the SURREAL [53] test set. As we can see, the fitted mesh has a similar error as our regression ones which uses the interpolation matrix $\hat{\mathbf{A}}$, which validates the accuracy of the estimated virtual markers.
B.4  Additional Qualitative Results
Figure 12 shows more qualitative comparisons with Pose2Mesh [7] on the SURREAL test set in which has diverse body shapes. The skeleton representation used in Pose2Mesh loses the body shape information so the method [7] can only recover mean shapes. For example, in Figure 12 (d) (e), the estimated meshes of Pose2Mesh tend to have the average body shape and fail to estimate the real body shape, regardless of whether the person is slim or stout. This is caused by the limited skeleton representation bottleneck so that the model learns a mean shape for the whole training dataset implicitly. In contrast, our approach with virtual marker representation generates more accurate mesh estimation results.
[…]
     Fitting      44.6      34.8      29.5
     Ours      44.7      36.9      28.9
Table 10: Results on SURREAL [53] test set when the mesh is obtained by fitting to the estimated virtual markers.
Figure 13 shows more qualitative comparisons with Pose2Mesh [7] and METRO [32] on the 3DPW test set. Pose2Mesh and METRO use the skeleton or all 3D vertices as intermediate representations, respectively. The estimated meshes are overlaid on the images according to the camera parameters. Pose2Mesh [7] has difficulty in estimating correct body pose and shapes when truncation occurs (a) or in complex postures (c). The results of METRO [32] have many artifacts where the estimated mesh is not smooth, and they also fail to align the image well. In contrast, our method estimates more accurate human poses and shapes and has smooth human mesh results. In addition, it is more robust to truncation and occlusion and aligns the image better.
Figure 11: Typical failure cases. (a) The right arm has inaccurate shape estimation due to the inaccurate virtual marker estimation around the arm when occluded. (b) Our method treats the lower arm of another person as its own due to occlusion. (c) Interpenetration around the right hand.
Figure 14 shows more quality results of our approach on the 3DPW [54], H3.6M [16], MPI-INF-3DHP [38], and COCO [34] datasets. Figure 10 shows more qualitative results on Internet images with challenging cases, such as extreme body shapes or complex poses. Our method generalizes well on the natural scenes. Figure 11 shows typical failure cases, including inaccurate shape estimation and interpenetration, which are mainly caused by inaccurate 3D virtual marker estimation when occlusion occurs. But as expected, the rest body parts are barely affected due to the local and sparse property of the virtual marker.
Figure 12: Qualitative comparison between our method and Pose2Mesh [7] on SURREAL test set [53]. Our approach generates more accurate mesh estimation results on images of diverse body shapes.
Figure 13: Qualitative comparison between our method and Pose2Mesh [7], METRO [32] on 3DPW test set [54]. Our approach is more robust to occlusion and truncation and generates more accurate mesh estimation results that align images well.
Figure 14: Meshes estimated by our approach on images from the 3DPW [54] dataset (row 1-4), H3.6M [16] dataset (row 5), MPI-INF-3DHP [38] dataset (row 6), and COCO dataset (last 2 rows) [34].
C.  Human Subject Data
We use existing public datasets of human subjects in our experiments following their official licensing requirements. With proper usage, the proposed method could be beneficial to society (e.g. elderly fall detection).